what is the length of each of a square that has a perimeter of 56cm
Answer by stanbon(75887)

what is the length of each of the dimensions of a square that has a perimeter of 56cm
-----
Perimeter = 2(L + W)
--
Note L = W
-----------
56 = 2(L + L)
56 = 2*2L
Length = 14 cm
L = Width = 14 cm
======================
Cheers,
Stan H.
======================